### v1

# Motion-R1: Chain-of-Thought Reasoning and Reinforcement Learning for Human Motion Generation

Abstract
Recent advances in large language models, especially in natural language understanding and reasoning, have opened new possibilities for text-to-motion generation. Although existing approaches have made notable progress in semantic alignment and motion synthesis, they often rely on end-to-end mapping strategies that fail to capture deep linguistic structures and logical reasoning. Consequently, generated motions tend to lack controllability, consistency, and diversity. To address these limitations, we propose Motion-R1, a unified motion-language modeling framework that integrates a Chain-of-Thought mechanism. By explicitly decomposing complex textual instructions into logically structured action paths, Motion-R1 provides high-level semantic guidance for motion generation, significantly enhancing the model’s ability to interpret and execute multi-step, long-horizon, and compositionally rich commands. To train our model, we adopt Group Relative Policy Optimization, a reinforcement learning algorithm designed for large models, which leverages motion quality feedback to optimize reasoning chains and motion synthesis jointly. Extensive experiments across multiple benchmark datasets demonstrate that Motion-R1 achieves competitive or superior performance compared to state-of-the-art methods, particularly in scenarios requiring nuanced semantic understanding and long-term temporal coherence. The code, model and data will be publicly available.
1 Introduction
Text-to-Motion (T2M) generation guoBackMLPSimple2022a ; tevet2022human ; Guo_2022_CVPR has emerged as a fundamental task in human-machine interaction, enabling the synthesis of realistic human motions from natural language descriptions. Driven by rapid advancements in large language models (LLMs) achiam2023gpt ; li2023blip ; qi2024cocogesture , recent T2M approaches  jiangMotionGPTHumanMotion2023 ; wangMotionGPT2GeneralPurposeMotionLanguage2024 ; wang2024motionagent have made significant strides in generating high-fidelity motions that align with complex textual instructions. Despite this progress, current approaches predominantly rely on end-to-end supervised learning ahnText2ActionGenerativeAdversarial2017 ; hongAvatarCLIPZeroShotTextDriven2022 ; ahujaLanguage2PoseNaturalLanguage2019 , mapping textual inputs directly onto motion sequences. While straightforward, such methods often fail to capture deeper linguistic structures and complex logical relationships embedded in natural language. For example, a high-level instruction such as "a person makes a cup of coffee" implicitly involves a sequence of semantically connected sub-actions, such as reaching, grasping, pouring, stirring, and placing, which require temporal ordering and causal reasoning. End-to-end models often struggle to interpret and execute such compositional commands accurately, resulting in incoherent or oversimplified motions.
Figure 1: Motion-R1 results on Out-of-Distribution prompts. Left: Complex caption with multi-step reasoning. Right: Abstract caption requiring semantic understanding. Motion-R1 captures structure and intent in both cases.
Inspired by the recent success of Chain-of-Thought (CoT) weicot2022 prompting in enhancing reasoning capabilities of LLMs, we hypothesize that explicitly modeling intermediate reasoning steps can similarly benefit motion generation tasks. By decomposing high-level instructions into structured action plans, models can better capture temporal dependencies, hierarchical goals, and fine-grained semantic nuances embedded in language. Moreover, reinforcement learning (RL) kaelbling1996reinforcement , especially Group Relative Policy Optimization (GRPO) shao2024deepseekmathpushinglimitsmathematical , offers a natural framework for refining such reasoning and generation processes by optimizing them for downstream motion quality. As shown in Figure 1, combining CoT-style semantic decomposition with RL-based optimization can improve the model’s generalization to out-of-distribution inputs, enhance the interpretability of action sequences, and facilitate more precise control over complex motion synthesis.
To this end, we propose Motion-R1, a unified T2M framework that integrates a CoT mechanism. By combining structured reasoning with reward-driven optimization, Motion-R1 significantly enhances the model’s generalization ability, capability to explore diverse motion styles, and interpretability of generated motions.
Specifically, we propose the MotionCoT Data Engine, an automated pipeline that leverages LLMs to generate high-quality Chain-of-Thought (CoT) annotations for motion descriptions. By distilling the semantic planning abilities of LLMs into structured action guidance, it helps the model capture the underlying logic of human actions. This enriched supervision improves generalization in motion synthesis. Compared to manual annotation, our approach produces more diverse and consistent CoT data at lower cost, serving as both an effective training signal and an interpretable bridge between text and motion.
We extend GRPO to the motion generation domain, offering a lightweight and scalable alternative to conventional RL. GRPO requires no value networks (e.g., PPO) or human preference data (e.g., DPO rafailov2023direct , RLHF), and instead uses group-wise ranking to learn from quality differences in generated samples—enhancing both generalization and diversity. To improve semantic alignment, we design a multi-aspect reward tailored to motion synthesis and adopt supervised fine-tuning (SFT) as a cold-start strategy for stable RL training.
The contributions of our work are summarized as follows:
$\bullet$ We introduce Motion-R1, a reasoning-guided and reinforcement-optimized framework for text-to-motion generation. It enhances the model’s ability to interpret abstract and complex textual descriptions through CoT planning and improves generalization to out-of-distribution instructions.
$\bullet$ We develop MotionCoT Data Engine, an automated CoT annotation pipeline that leverages LLMs to generate structured motion planning paths. This enables effective distillation of LLMs’ reasoning capabilities and reduces reliance on costly manual labeling.
$\bullet$ We extend the GRPO algorithm to the motion generation domain, offering a simple yet effective reinforcement learning paradigm that requires no value networks or preference data. Additionally, we propose multi-aspect reward functions for motion synthesis to strengthen semantic grounding and training stability.
$\bullet$ Extensive experiments demonstrate that Motion-R1 achieves state-of-the-art performance on multiple benchmarks. Compared with MotionAgent wang2024motionagent , our method exhibits superior semantic understanding and motion synthesis capabilities, particularly for complex and out-of-distribution textual descriptions. Specifically, Motion-R1 outperforms MotionAgent by achieving significant improvements of $\mathbf{1.19}\%$ in motion diversity and $\mathbf{2.12}\%$ in R-precision.
2 Related Work
2.1 Multimodal Large Language Models
Recently, multimodal large language models (MLLMs) wu2023multimodal ; koh2023generating ; tang2025video have shown impressive performance in various tasks, including text generation, reasoning, and even multimodal tasks. For example, models like GPT-4 achiam2023gpt and BLIP-2 li2023blip have demonstrated the ability to understand and generate text and images. While recent MLLMs have significantly improved their perception and generation capabilities across diverse modalities, many complex real-world tasks demand not only understanding but also reasoning banerjeeWeaklySupervisedRelative2021 ; xiongTEILPTimePrediction2024 . Consequently, enhancing the reasoning abilities of MLLMs has become an important research direction.
Building on the success of CoT  weicot2022 prompting in NLP, researchers have explored extending CoT-style reasoning to multimodal settings. For instance, models like LLaVA-CoT xuLLaVACoTLetVision2025 introduce a novel vision-language model capable of performing autonomous, structured, and multistage reasoning by decomposing complex questions into four stages: summary, caption, reasoning, and conclusion. While structured reasoning improves the interpretability and controllability of vision-language models, recent studies tan2025reason ; pan2025medvlm ; liu2025seg further suggest that reasoning capabilities can be learned and optimized via feedback-driven training. DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 introduces GRPO, a reinforcement learning algorithm that trains LLMs to produce more accurate and coherent responses, leveraging reward signals from downstream tasks. Although this approach has shown promise in improving the reasoning capabilities of MLLMs huangVisionR1IncentivizingReasoning2025 , its application in the context of 3D human motion generation remains unexplored. This idea inspires our approach, which extends GRPO to the domain of motion generation, using action quality as feedback to refine both the thought process and the generated motion sequence.
2.2 Text-guided 3D Human Motion Generation
Text-guided 3D human motion generation has become a key research focus in recent years. Early methods ahnText2ActionGenerativeAdversarial2017 ; ghoshSynthesisCompositionalAnimations2023 ; qi2024weakly primarily relied on generative adversarial networks goodfellow2014generative ; wang2024hallo3d to synthesize human motion from text descriptions. These methods often struggled with issues such as limited diversity.
To address these challenges, diffusion models ho2020denoising ; li2024era3d ; li2025pshuman have gained popularity in this field. Notable works zhouUDEUnifiedDriving2022 ; kimFLAMEFreeformLanguagebased2023 ; dabralMoFusionFrameworkDenoisingDiffusionbased2023 include MDM (tevet2022human, ), which first applies diffusion modeling to motion synthesis; MotionDiffuse (zhang2023motiondiffuse, ) improves temporal consistency; MotionChain (jiang2023motionchain, ) explores fine-grained structure and staged modelling; and MoMask (guo2023momask, ) employs masking strategies for improved training efficiency. However, these methods often rely on complex architectures and extensive training data and are limited by their motion initialization.
In parallel, VQ-VAE-based methods Guo_2022_CVPR ; hongAvatarCLIPZeroShotTextDriven2022 ; petrovichTEMOSGeneratingDiverse2022 ; guoTM2TStochasticTokenized2022 ; athanasiouTEACHTemporalAction2022 discretize motion representations to facilitate efficient sequence modeling and better alignment with language, providing a more interpretable latent space. Building on this, LLMs have been integrated into the generation process to enhance the quality and control of generated motions. For instance, T2M-GPT (zhangT2MGPTGeneratingHuman2023, ) combines VQ-VAE with GPT in a two-stage framework for high-quality generation. Subsequent works such as MotionDiffuse zhangMotionDiffuseTextDrivenHuman2022 , MotionCLIP (tevet2022motionclip, ), and MotionGPT (jiangMotionGPTHumanMotion2023, ) integrate diffusion modeling, contrastive learning, and Transformer architectures to enhance fidelity and semantic consistency further. MotionAgent (wang2024motionagent, ) further incorporates semantic planning for robust generalization in complex task scenarios. Meanwhile, with the development of reinforcement learning, MotionRL liuMotionRLAlignTexttoMotion2024 leverages PPO schulman2017proximal to improve the generation based on human preferences prior knowledge.
However, despite these advances, current T2M methods largely adopt an end-to-end mapping mode, lacking explicit reasoning capabilities over complex linguistic instructions. Our approach aims to address this gap by introducing a CoT-guided and reward-optimized framework for 3D human motion generation, enhancing the interpretability and diversity of the generated motions.
3 Method
Figure 2: Motion-R1 pipeline. The framework consists of two stages: (1) MotionCoT Data Engine uses DeepSeek-R1 to generate CoT-style motion planning traces in <think>, <output>, and <Motion> format to fine-tune an LLM; (2) GRPO-based training ranks grouped outputs by format, motion, and semantic rewards to optimize the LLM via reinforcement learning.
3.1 Motion-R1 Framework
Motion-R1 comprises two core components: a pre-trained motion tokenizer and an LLM equipped with action-oriented reasoning capabilities. The motion tokenizer discretizes continuous motion sequences into motion tokens and reconstructs them back into smooth, coherent trajectories. On the other hand, the LLM is designed to perform structured reasoning over natural language instructions, enabling it to decompose complex action descriptions into finer-grained, logically ordered sub-actions — what we refer to as a reasoning-based motion breakdown. Based on this structured interpretation, the model generates high-quality motion token sequences that faithfully reflect the intended behavior.
As shown in Figure 2, our framework consists of two training stages. In the first stage, we construct a motion-language dataset with CoT supervision weicot2022 by prompting DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 . This dataset, referred to as the MotionCoT Data Engine, enables cold-start supervised tuning of the LLM to produce structured reasoning process in the <think>, <output>, and <Motion> format.
In the second stage, we refine the LLM via reinforcement learning using GRPO shao2024deepseekmathpushinglimitsmathematical . GRPO operates by sampling $G$ generations per prompt and ranking them within a group. For each sampled output, a set of task-specific rewards $R_{n}$ is computed, including format correctness, motion similarity, and semantic alignment. Based on these rankings, relative advantages $A_{n}$ are estimated to update the policy without requiring an explicit value function. These feedback signals guide the LLM to generate more accurate and realistic motion sequences conditioned on textual instructions.
3.2 Motion Tokenizer
To integrate motion data, which differs significantly from natural language in structure and modality, into the LLM framework, we adopt a VQ-VAE architecture as our motion tokenizer. This approach has been widely adopted in  zhangT2MGPTGeneratingHuman2023 ; jiangMotionGPTHumanMotion2023 ; guo2023momask ; wang2024motionagent and proven effective for 3D human motion data modeling.
The motion tokenizer comprises an encoder $E$ and a decoder $D$. Given an input motion sequence $\mathbf{m}_{1:T}\in\mathbb{R}^{T\times D}$, where $T$ denotes the number of frames and $D$ is the number of dimensions of each frame, the encoder $E$ maps the sequence to a latent representation $\mathbf{z}_{1:(T/l)}\in\mathbb{R}^{(T/l)\times d}$, where $l$ is the temporal downsampling rate and $d$ is the number of latent dimensions. Each latent vector $\mathbf{z_{i}}$ is then quantized using a learnable codebook $\mathbf{C}=\{\mathbf{c}_{n}\}_{n=1}^{N}$, where $N$ is the codebook size and $\mathbf{c}_{n}\in\mathbb{R}^{d}$ represents a discrete motion token. The quantization selects the nearest code vector to each embedding:
$\hat{\mathbf{z}}_{i}=\arg\min_{\mathbf{c}_{n}\in\mathbf{C}}\|\mathbf{z}_{i}-\mathbf{c}_{n}\|_{2}$ (1)
The original motion sequence is reconstructed as $\hat{\mathbf{m}}_{1:T}=D(\hat{\mathbf{z}}_{1:(T/l)})$. Following zhangT2MGPTGeneratingHuman2023 , we train the VQ-VAE model using a composite objective that includes a reconstruction loss $L_{\text{reconstruct}}$, a codebook commitment loss $L_{\text{commit}}$, and an embedding loss $L_{\text{embed}}$:
$L_{\text{vq}}=L_{\text{reconstruct}}+L_{\text{commit}}+L_{\text{embed}}$ (2)
Here, $L_{\text{reconstruct}}$ includes a smoothed L1 loss with velocity regularization to improve generation quality. To ensure stable and efficient training, we follow  zhangT2MGPTGeneratingHuman2023 to incorporate exponential moving average (EMA) updates for the codebook along with codebook reset strategies.
Figure 3: MotionCoT Data Example. Given the prompt “A person does Tai chi.” the LLM generates a step-by-step CoT reasoning trace (<think>) and a structured action plan <output>, covering stance, arm movement, weight transfer, and hand positioning.
3.3 MotionCoT Data Engine
We present MotionCoT Data Engine, an automated framework for constructing CoT weicot2022 supervision tailored to the text-to-motion generation task. This engine leverages the general reasoning and semantic planning capabilities of LLM to transform free-form motion descriptions into structured, high-quality CoT action plans, which serve as intermediate guidance between language and motion.
The data engine begins by augmenting existing motion-language datasets through prompt-based querying. We design a comprehensive prompt that instructs the LLM to reason through the detailed logic and temporal structure. This prompt includes clear task instructions, output format constraints, and several in-context examples to guide the model in producing structured, step-by-step CoT reasoning. The detailed prompt is provided in the supplementary material.
The CoT outputs generated by this process are then passed through an automated quality control pipeline. Specifically, we employ DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 to evaluate each generated CoT for relevance, conciseness, and logical coherence. Samples exhibiting redundancy, overthinking, or excessive verbosity are filtered out and regenerated iteratively until they meet our quality standards. As shown in Figure 3, we give an example of the CoT generation process. The LLM is prompted to generate a CoT reasoning process for the task of “A person does Tai chi”. The model first identifies the main action, “Tai chi”, and then breaks it down into sub-actions such as “stand up”, “arm movements”, “weight displacement”, and “hand positioning”. Each sub-action is further elaborated with specific details, such as the direction of movement and the body parts involved.
Once validated, each high-quality CoT trace is paired with its original motion description and corresponding action sequence, forming triplets of the form (description, CoT, motion). These triplets constitute a refined dataset that provides explicit, interpretable reasoning supervision for downstream training.
By distilling structured action plans from LLMs, the MotionCoT Data Engine enables the model to learn fine-grained and context-aware semantic decomposition from natural language. This significantly enhances the model’s controllability, interpretability, and generalization capability in complex motion generation tasks while dramatically reducing the cost of manual annotation.
3.4 Training Strategy
3.4.1 Cold-Start Training with CoT
Inspired by the success of DeepSeek-R1-Zero deepSeekR1IncentivizingReasoning2025 in solving mathematical reasoning tasks via pure RL, we initially explore end-to-end RL training for our model, aiming to jointly induce CoT reasoning and motion token generation solely from reward signals. However, this approach proves highly unstable in our setting: the model frequently fails to generate coherent CoT sequences and, more critically, produces motion tokens that are either semantically misaligned or syntactically invalid.
These challenges may be partially explained by two fundamental differences between our task and the conventional domains of symbolic reasoning. First, unlike the short-form outputs (e.g., numerical answers) in math problems, motion generation requires long, structured sequences of discrete tokens. Second, these motion tokens are newly introduced symbols in the model’s vocabulary, whose embeddings are insufficiently trained to bridge the modality gap between natural language and motion representations. As a result, the model struggles to interpret and generate them correctly.
To address these issues, we adopt a cold-start strategy based on supervised fine-tuning. Leveraging the curated dataset, comprising aligned triplets of task descriptions, CoT reasoning process, and motion sequences, we fine-tune the language model to bootstrap its ability to generate both structured CoT explanations and valid motion outputs. This supervised initialization constrains the model to learn the expected format and semantics of both reasoning and motion, providing a stable and effective foundation for subsequent policy optimization via RL.
3.4.2 GRPO-based Reinforcement Learning
To bridge the gap between natural language semantics and motion synthesis, we formulate text-to-motion generation as an RL problem and adopt GRPO shao2024deepseekmathpushinglimitsmathematical to optimize motion generation policies without relying on explicit reward functions or human-labeled preferences.
Unlike traditional RL approaches such as PPO schulman2017proximal or DPO rafailov2023direct , which require value networks or pairwise preference data, GRPO is a lightweight and scalable alternative. It leverages intra-group ranking feedback to optimize the generation policy by encouraging better samples within each group. This eliminates the need for training an additional critic model and reduces the overall computational overhead.
For each text prompt $q$, a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ are sampled from the old policy model $\pi_{\text{old}}$. Each output is assigned a scalar reward, yielding a reward vector $\mathbf{r}=\{r_{1},r_{2},\dots,r_{G}\}$, computed by task-specific reward functions that evaluate the quality of each output. GRPO then updates the policy model by maximizing the following clipped objective:
$\mathcal{J}_{\text{GRPO}}(\theta)=$ (3)
$\mathbb{E}_{c}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\text{old}}(o_{i}|q)}\hat{A}_{i},\;\text{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\text{old}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\right)\hat{A}_{i}\right)-\beta\cdot D_{\text{KL}}(\pi_{\theta}\,\|\,\pi_{\text{ref}})\right],$
where $\varepsilon$ and $\beta$ are hyperparameters controlling the clipping range and KL regularization strength, respectively. $\pi_{\text{ref}}$ denotes the reference policy model. The normalized advantage term is given by $\hat{A}_{i}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$. This formulation enables stable and efficient policy optimization, incorporating semantic- and motion-aware reward shaping tailored specifically for the text-to-motion generation task.
To better guide policy learning and improve alignment between textual prompts and generated motions, we introduce a multi-reward mechanism.
Format Reward. To ensure that the content generated by the model has a resolvable structure, we introduce Format Reward $r_{\text{format}}$. This reward detects through regularization expressions whether the generated results strictly follow the predefined format: <think>{CoT}</think><Motion>{Motion tokens}</Motion>. Here, curly braces and their enclosed content denote placeholders, representing the generated CoT and the corresponding motion tokens, respectively. If the generated content exactly matches this required format, we assign a reward of 1. Otherwise, the reward is set to 0.
Motion Similarity Reward. To encourage temporal and spatial coherence between the generated motion $\hat{\mathbf{m}}$ and the ground-truth motion $\mathbf{m}$, we adopt a motion similarity reward $r_{\text{motion}}$. Specifically, we employ a pre-trained motion encoder $f_{\text{motion}}$, as introduced in Guo_2022_CVPR , to obtain feature representations of both motions. The reward is computed as the cosine similarity between their embeddings:
$r_{\text{motion}}=\frac{f_{\text{motion}}(\hat{\mathbf{m}})\cdot f_{\text{motion}}(\mathbf{m})}{\|f_{\text{motion}}(\hat{\mathbf{m}})\|_{2}\cdot\|f_{\text{motion}}(\mathbf{m})\|_{2}}$ (4)
Semantic Similarity Reward. To ensure that the generated motion semantically corresponds to the input textual description $T$, we introduce the semantic similarity reward $r_{\text{semantic}}$. This reward measures the alignment between the motion embedding and the language embedding in a shared latent space, both extracted using pre-trained encoders $f_{\text{motion}}$ and $f_{\text{text}}$, respectively (also from Guo_2022_CVPR ). The reward is defined as:
$r_{\text{semantic}}=\frac{f_{\text{motion}}(\hat{\mathbf{m}})\cdot f_{\text{text}}(T)}{\|f_{\text{motion}}(\hat{\mathbf{m}})\|_{2}\cdot\|f_{\text{text}}(T)\|_{2}}$ (5)
These reward signals are not directly used in scalar form but are reflected through groupwise preference ranking in GRPO.
4 Experiments
4.1 Experimental Settings
Datasets We evaluate our method on two widely-used benchmarks: HumanML3D Guo_2022_CVPR and KIT-ML plappertKITMotionLanguageDataset2016 . HumanML3D contains $14,616$ motions and $44,970$ textual descriptions, sourced AMASS AMASS2019 and HumanAct12 guo2020action2motion . KIT-ML includes $3,911$ motion clips and $6,278$ descriptions, derived from KIT kit2015 and CMU CarnegieMellonUniversity datasets. Both datasets follow standard train/val/test splits.
Table 1: Quantitative results of Motion-R1 on HumanML3D. The evaluations are conducted 20 times to obtain a $95\%$ confidence interval. Best results are highlighted in bold and the second best in underline.
[…]
Top 1 Top 2 Top 3
MDM (tevet2022human, ) $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $\textbf{2.799}^{\pm.072}$
MLD (chenExecutingYourCommands2023, ) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
MotionDiffuse (zhang2023motiondiffuse, ) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
T2M (Guo_2022_CVPR, ) $0.457^{\pm.002}$ $0.559^{\pm.007}$ $0.740^{\pm.003}$ $1.067^{\pm.002}$ $3.340^{\pm.008}$ $9.188^{\pm.002}$ $2.090^{\pm.083}$
TM2T (guoTM2TStochasticTokenized2022, ) $0.424^{\pm.003}$ $0.618^{\pm.003}$ $0.729^{\pm.002}$ $1.501^{\pm.017}$ $3.467^{\pm.011}$ $8.589^{\pm.076}$ $\underline{2.424}^{\pm.093}$
T2M-GPT (zhangT2MGPTGeneratingHuman2023, ) $0.491^{\pm.003}$ $0.680^{\pm.003}$ $0.775^{\pm.002}$ $\underline{0.116}^{\pm.004}$ $3.118^{\pm.011}$ $9.761^{\pm.081}$ $1.856^{\pm.011}$
MotionGPT (jiangMotionGPTHumanMotion2023, ) $0.492^{\pm.003}$ $0.681^{\pm.003}$ $0.778^{\pm.002}$ $0.232^{\pm.008}$ $3.096^{\pm.008}$ $9.528^{\pm.071}$ $2.008^{\pm.084}$
MoMask guo2023momask $\textbf{0.521}^{\pm.002}$ $\underline{0.713}^{\pm.002}$ $\underline{0.807}^{\pm.002}$ $\textbf{0.045}^{\pm.002}$ $\underline{2.958}^{\pm.008}$ ${9.620}^{\pm.064}$ $1.241^{\pm.040}$
MotionChain (jiang2023motionchain, ) $0.504^{\pm.003}$ $0.617^{\pm.002}$ $0.790^{\pm.003}$ $0.248^{\pm.009}$ $3.033^{\pm.010}$ $9.470^{\pm.075}$ $1.727^{\pm.014}$
MotionLLM wang2024motionagent $0.515^{\pm.004}$ $0.691^{\pm.003}$ $0.801^{\pm.004}$ $0.230^{\pm.009}$ $2.967^{\pm.020}$ $9.908^{\pm.102}$ $2.142^{\pm.014}$
MotionGPT-2 wangMotionGPT2GeneralPurposeMotionLanguage2024 $0.496^{\pm.002}$ $0.691^{\pm.003}$ $0.782^{\pm.004}$ $0.191^{\pm.004}$ $3.080^{\pm.013}$ $9.860^{\pm.026}$ $2.137^{\pm.022}$
Motion-R1 $\underline{0.515}^{\pm.003}$ $\textbf{0.719}^{\pm.002}$ $\textbf{0.818}^{\pm.002}$ $0.201^{\pm.004}$ $\textbf{2.854}^{\pm.010}$ $\textbf{10.026}^{\pm.075}$ $2.317^{\pm.105}$
Evaluation Metrics Following standard protocols Guo_2022_CVPR ; zhangT2MGPTGeneratingHuman2023 ; wang2024motionagent , we report R-Precision@1/2/3, FID, Diversity, MM-Dist, and MModality. These metrics respectively evaluate retrieval accuracy, distributional realism, sample diversity, text-motion alignment, and one-to-many generation ability. Note that more details about the evaluation metrics can be found in the supplementary material.
4.2 Implementation Details
For constructing the MotionCoT Data Engine, we utilize DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 (API Version: 250120) to generate detailed CoT explanations corresponding to complex motion descriptions. These structured reasoning traces serve as intermediate semantic representations that bridge textual instructions and action planning. In Motion-R1, we adopt Qwen-2.5-3B-Instruct qwen25 as the backbone language model for generating CoT sequences and guiding motion synthesis. Qwen-2.5 is selected for its strong performance in multi-step reasoning and compact model size, which ensures efficient training within our RL pipeline. The generated CoTs are subsequently used as inputs to the motion generator and as targets for reward-based optimization via GRPO. All of our experiments are conducted on NVIDIA H20 GPUs.
4.3 Quantitative Evaluation
We evaluate the performance of Motion-R1 on HumanML3D and KIT-ML datasets, comparing it with state-of-the-art methods, ranging from VAE-based to diffusion-based models. For the diffusion-based model, we select MDM tevet2022human , MLD chenExecutingYourCommands2023 , MotionDiffuse zhang2023motiondiffuse . For the VAE-based model, we choose T2M Guo_2022_CVPR , TM2T guoTM2TStochasticTokenized2022 , T2M-GPT zhangT2MGPTGeneratingHuman2023 , MotionGPT jiangMotionGPTHumanMotion2023 , MoMask guo2023momask , MotionChain jiang2023motionchain , MotionLLM wang2024motionagent and MotionGPT-2 wangMotionGPT2GeneralPurposeMotionLanguage2024 .
We utilize the same evaluation settings as in previous works Guo_2022_CVPR ; wang2024motionagent , ensuring a fair comparison. For each method, we report the average results across 20 times (except MModality 5 times) to obtain a $95\%$ confidence interval. The quantitative results on HumanML3D are presented in Table 1. As shown in Table 1, Motion-R1 achieves state-of-the-art performance across multiple evaluation metrics on the HumanML3D dataset. In terms of R-Precision, Motion-R1 attains an R-Precision@1 of $0.515$, which is on par with the best-performing method (MoMask guo2023momask ), while achieving the highest R-Precision@2 ($0.719$) and R-Precision@3 ($0.818$), demonstrating superior semantic alignment and reasoning capabilities in motion-to-text retrieval. In terms of motion quality, Motion-R1 achieves a low FID of $0.201$, which is competitive with recent strong baselines such as MotionGPT-2 ($0.191$) and MotionGPT ($0.232$), indicating high realism and fidelity of the generated motions. Our method outperforms all baselines in MM-Dist ($2.854$), highlighting better alignment in the shared embedding space of motion and language. Moreover, Motion-R1 significantly outperforms all competing methods on the Diversity metric, reaching a score of $10.026$, suggesting that the model effectively captures the one-to-many nature of text-to-motion mappings. On the MModality metric, our approach achieves a strong score of $2.317$, which is competitive with other top methods, further validating the richness of generated motions conditioned on the same textual input.
Figure 4: Visualization comparisons with MotionLLM wang2024motionagent on in-distribution and out-of-distribution prompts.
4.4 Qualitative Analysis and Visualization
To complement the quantitative evaluation, we provide qualitative comparisons between Motion-R1 and existing state-of-the-art methods. We divide this section into two parts: (1) in-distribution prompts from standard benchmarks, and (2) out-of-distribution instructions that require compositional reasoning and generalization. These visualizations highlight the advantages of Motion-R1 in producing semantically coherent, diverse, and controllable motion sequences.
In-Distribution Visualization. We first visualize motion generation results on descriptions drawn from the HumanML3D and KIT-ML test sets. These examples cover various action types, including short actions and multi-step descriptions. We compare Motion-R1 against representative baselines such as MoMask and MotionGPT. As shown in Figure 4 (left), Motion-R1 produces smooth, well-structured sequences for simple and multi-step instructions. For example, given “A person walks in a circle.” our method generates a continuous loop with clear orientation change and natural timing, while MotionLLM wang2024motionagent fails to complete a full circle or exhibits abrupt terminations. This illustrates Motion-R1’s better alignment with spatial and temporal semantics, even in standard settings.
Table 2: Ablation study on HumanML3D. CoT, $R_{\text{sem}}$, and $R_{\text{motion}}$ denote the Chain-of-Thought decomposition, the semantic similarity reward, and the motion similarity reward, respectively. The evaluations are conducted 20 times to obtain a $95\%$ confidence interval. Best results are highlighted in bold and the second best in underline.
[…]
Top 1 Top 2 Top 3
$\checkmark$ ${0.340}^{\pm.002}$ $0.503^{\pm.002}$ $0.603^{\pm.002}$ $0.530^{\pm.015}$ $4.216^{\pm.015}$ $9.696^{\pm.090}$ $\textbf{4.762}^{\pm.249}$
$\checkmark$ $\checkmark$ $0.482^{\pm.002}$ $\underline{0.690}^{\pm.003}$ $\underline{0.799}^{\pm.001}$ $0.297^{\pm.004}$ $2.963^{\pm.004}$ $9.537^{\pm.021}$ ${2.317}^{\pm.015}$
$\checkmark$ $\checkmark$ $\underline{0.483}^{\pm.002}$ $\underline{0.690}^{\pm.002}$ $\underline{0.799}^{\pm.002}$ $\underline{0.281}^{\pm.008}$ $\underline{2.947}^{\pm.007}$ $9.848^{\pm.082}$ $1.903^{\pm.276}$
$\checkmark$ ${0.322}^{\pm.002}$ ${0.497}^{\pm.003}$ ${0.610}^{\pm.002}$ $1.705^{\pm.034}$ ${4.206}^{\pm.010}$ $\underline{9.914}^{\pm.057}$ $\underline{4.194}^{\pm.192}$
$\checkmark$ ${0.365}^{\pm.002}$ ${0.526}^{\pm.001}$ ${0.627}^{\pm.003}$ $0.766^{\pm.021}$ ${4.136}^{\pm.005}$ ${9.711}^{\pm.041}$ $4.067^{\pm.200}$
$\checkmark$ $\checkmark$ ${0.365}^{\pm.002}$ ${0.539}^{\pm.003}$ ${0.644}^{\pm.002}$ $1.656^{\pm.015}$ ${4.024}^{\pm.012}$ ${9.720}^{\pm.024}$ $4.008^{\pm.119}$
$\checkmark$ $\checkmark$ $\checkmark$ $\textbf{0.515}^{\pm.003}$ $\textbf{0.719}^{\pm.002}$ $\textbf{0.818}^{\pm.002}$ $\textbf{0.201}^{\pm.004}$ $\textbf{2.854}^{\pm.010}$ $\textbf{10.026}^{\pm.075}$ $2.317^{\pm.105}$
Out-of-Distribution Visualization. To evaluate generalization beyond the training distribution, we present qualitative comparisons under two types of out-of-distribution captions, as shown in Figure  4 (middle and right).
The first prompt describes a complex, multi-stage action with implicit causal logic: “After hearing a loud noise, a person turned around, stepped back cautiously with hands raised defensively, and then slowly approached.” Motion-R1 captures this chain of reactions in a temporally coherent manner, clearly separating reaction, retreat, and re-approach. In contrast, MotionLLM either merges steps or lacks defensive gestures, resulting in an incomplete or disordered motion sequence.
The second prompt—“A person is serving in badminton”—involves abstract semantics and task-specific dynamics. Motion-R1 synthesizes a plausible serve-like motion, including arm lifting, body shift, and a forward striking gesture, resembling real badminton actions. MotionLLM, however, produces repetitive or generic motions lacking physical intent or domain alignment.
These results demonstrate that Motion-R1 generalizes well to unseen instructions by leveraging explicit CoT-style reasoning and planning, while baselines such as MotionLLM struggle with long-horizon dependencies or abstract concepts.
4.5 Ablation Study
To assess the contribution of each component in our framework, we conduct an ablation study on the HumanML3D dataset, as shown in Table 2. Specifically, we evaluate the impact of three key components: CoT, the semantic similarity reward ($R_{\text{sem}}$), and the motion similarity reward ($R_{\text{motion}}$).
All ablated variants achieve comparable performance in most metrics, reflecting the robustness of our design. Notably, enabling CoT alone results in the poorest performance, with significantly lower R-Precision scores (Top-1: $0.340{\pm}.002$) and higher FID ($0.530{\pm}.015$) and MM-Dist ($4.216{\pm}.015$), indicating insufficient alignment between generated motions and the input descriptions. Adding either $R_{\text{sem}}$ or $R_{\text{motion}}$ to CoT yields a substantial boost, particularly in R-Precision Top-1 (increasing to $0.483{\pm}.002$) and FID (dropping to $0.281{\pm}.008$), suggesting that semantic and motion-level feedback are both crucial for enhancing generation quality.
[…]
5 Conclusion and Limitations
Conclusion   In this work, we presented Motion-R1, a unified framework for text-to-motion generation that combines Chain-of-Thought reasoning and reinforcement learning. Our approach decomposes complex language into structured action plans and optimizes motion synthesis using GRPO, improving controllability, diversity, and generalization. To support reasoning-based training, we introduced the MotionCoT Data Engine for automatic CoT-style annotation.
Limitations   While Motion-R1 shows promising results, it still has limitations. The CoT decomposition is generated by prompting general-purpose LLMs, which may introduce noise or suboptimal planning in ambiguous instructions. Additionally, although GRPO simplifies the RL pipeline by removing the need for explicit value functions or preference data, it still requires careful design of multi-component rewards. Future work will explore adaptive reward learning and interactive feedback to further enhance generation quality.